Slight burning smell in the Music Building: bricklayers are using a saw in their activities, which is located near the air intake for the building. The smell will be intermittent as the outside air-dampers open and close throughout the day to provide the building with oxygen, and pull out CO2. There is nothing to be alarmed about.
Smith construction is ongoing - crews are putting final touches on 1st floor North rooms, and preparations to pour concrete in Smith 13 is underway.
Materials removal at Smith is incredibly loud. Students are encouraged to use the MB to practice today! Hearing protection is advised.
Practice room cleanups - equipment is being redistributed and rooms are being swept out.
MB 0303 - issue with Ethernet cable in the wall. The connection is not stable, and may need to be unplugged and re-plugged for the signal to appear on the monitor. ELMO had goo spilled on it and the HDMI connection is loosening up. Please use care with this unit.
MB 0304 - functioning correctly.
MB 0326 - functioning correctly.
MB 0328 - VERY dirty with coffee spills on the inside of the cabinet. Desk has been cleaned. Please folks - this is YOUR equipment.
MB 0359 - volume knob is not functioning properly and is functioning intermittently. Projector will remain with poor color until May when the unit is slated for replacement.
TIP: if the volume knob is malfunctioning, turn up your device volume and use the main volume on the amplifier in the rack unit, not the podium.